What is the volume and area of a sphere with a radius of 10cm
Answer by addingup(3677)

Volume: 4/3*Pi*r^3
= 4/3*Pi*10^3
= 4/3*Pi*1000
= 4/3*3141.6
= 4188.80 cm^3 this is the volume
---------------------------------------
Area of the sphere: I interpret you to say surface area, so I'll give you the area of the outside of the sphere, the surface area. IF you want the total area (inside and out) just double it.
----------------------------------------
Area: 4*Pi*r^2
= 12.57*100
= 1257 cm^2 this is the surface area.